OpenStudy (anonymous):

ok so there's a GP 1 + (1+r) + (1+r)^2 + .... a = 1 and r(common ratio) = (1+r) ??

OpenStudy (zugzwang):

Seems like it... was there a further question?

OpenStudy (blurbendy):

(1 + r)^2 / (1 + r) = 1 + r (1 + r ) / 1 = 1 + r So, yes.

OpenStudy (anonymous):

well i had to find the sum, so i got r + 2, but i just wanted to check that I was going in the right direction.

OpenStudy (zugzwang):

This is only true if |1 + r| < 1 Let's assume it's positive, for the sake of simplicity... 0 < 1 + r < 1 -1 < r < 0 so the sum of the infinite geometric series is given by \[\frac{a}{1-R}\] Where a is the first term and R is the common ratio. That gives \[\frac{1}{1-(1+r)}\]

OpenStudy (anonymous):

ok, now in the next question, i don't know if its a continuation or not, so help :P

OpenStudy (anonymous):

On 1st Jan 2000, a man borrowed Rs. A at a fixed compound interest rate of a%. He pays back Rs. P on 1st Jan 2001, and then again Rs. P on 1st Jan 2002. Let a/100 = r Show that the amount still owing in Rs., just after this second payment, is A(1+r)^2 - P(1+r) - P.

OpenStudy (zugzwang):

I'm afraid this is beyond my scope :(

OpenStudy (anonymous):

awww.